How many liters of 20% alcohol and 75% alcohol must be mixed to obtain 32L of 40% alcohol solution?
Answer by stanbon(75887)

How many liters of 20% alcohol and 75% alcohol must be mixed to obtain 32L of 40% alcohol solution?
----
Equation:
alc + alc = alc
0.20x + 0.75(32-x) = 0.40*32
----
20x + 75*32 - 75x = 40*32
----
-55x = -35*32
x = (7/11)32
----
x = 20.36 liters (amt. of 20% solution needed)
-----
32-x = 32-20.36 = 11.64 liters (amt. of 75% solution needed)
===================
Cheers,
Stan H.
===================